Find the number b such that the line y=b divides the region bounded by the curves y=x2 and y=4 into two regions with equal area.


1
Expert's answer
2021-09-13T00:04:27-0400

Find the area of the region bounded by the curves y=x2 and y=4 


"x^2=4"

"x_1=-2, x_2=2"

"A_1=\displaystyle\int_{-2}^{2}(4-x^2)dx=\big[4x-\dfrac{x^3}{3}\big]\begin{matrix}
 2 \\
 -2
\end{matrix}"

"=8-\dfrac{8}{3}-(-8+\dfrac{8}{3})=\dfrac{32}{3} ({units}^2)"

Find the area of the region bounded by the curves y=x2 and y=b 


"x^2=b, b>0"

"x_1=-\sqrt{b}, x_2=\sqrt{b}"

"A_2=\displaystyle\int_{-\sqrt{b}}^{\sqrt{b}}(b-x^2)dx=\big[bx-\dfrac{x^3}{3}\big]\begin{matrix}
 \sqrt{b} \\
 -\sqrt{b}
\end{matrix}"

"=b\sqrt{b}-\dfrac{b\sqrt{b}}{3}-(-b\sqrt{b}+\dfrac{b\sqrt{b}}{3})=\dfrac{4b\sqrt{b}}{3} ({units}^2)"

The line y=b divides the region bounded by the curves y=x2 and y=4 into two regions with equal area


"A_1=2A_2"

"\dfrac{32}{3}=2(\dfrac{4b\sqrt{b}}{3})"


"b\sqrt{b}=4"

"b^3=16"

"b=2\sqrt[3]{2}"